I need help on trig identities.
Verify the following equation:
cosx(cosx-secx)=-sin^2x
So far I used the foil method to get cos^2x-cosxsecx but I'm not sure if that is right

Answer by lwsshak3(11628)

Verify the following equation:
cosx(cosx-secx)=-sin^2x
start with left side
cosx(cosx-secx)
=cosx(cosx-1/cosx)
=cos^2x-1
=-(1-cos^2x)
=-sin^2x
verified:
left side=right side